OpenStudy (anonymous):

Find the value of a: 3a = -1/4 a = 1/12 a = -1/12 a = -12 a = 12

OpenStudy (idealist10):

3a=-1/4 a=(-1/4)/3 a=(-1/4)(1/3) a=-1/12 Answer: -1/12

OpenStudy (anonymous):

3a=-1/4 divide by 3 a=-1/12